OpenStudy (stuck-help):

factor completely 48y^2+82y-70 metal+fan

OpenStudy (stuck-help):

show you work please

OpenStudy (welshfella):

first make it a bit easier by dividing through by 2 48y^2+82y-70 = 2(24y^2 + 41y - 35)

OpenStudy (welshfella):

now you need to factor the part in the parentheses find the prime factors of 24 and -35 then you need to combine them in such a way as to get the middle term in the expression (+41)

OpenStudy (welshfella):

24 = 2 * 2* 2 * 3 -35 = +/- 5 * +/- 7

OpenStudy (welshfella):

8*7 - 3*5 will give us + 41 we need 3x and 8x to be the first part in brackets to give the 24x^2 so the factors will look like (8x + a )(3x + b) except that to get the -35 one of the signs will be -

OpenStudy (welshfella):

can you work out what a and b will be?

OpenStudy (welshfella):

these are a bit tricky and need some practice to become proficient

OpenStudy (stuck-help):

im done with this question but thanks anyway

OpenStudy (welshfella):

yw